Two aqueous ethanol solutions with different concentrations were mixed. Mixing resulted in a solution with a concentration of ________ M. Solution A, 9.54 kg, has a concentration of 28.4%. Solution B, 5.27 L, has a concentration of 80.0 proof. Fill in the blank. ρethanol = 789 kg/m3


1
Expert's answer
2021-10-05T23:57:57-0400

9.54×0.284= 2.71kg

5.27×0.8= 4.21L

2.71/0.789= 3.435

3.435/4.21

= 0.816mol/L